# Long-lived quasinormal modes, grey-body factors and absorption cross-section of the black hole immersed in the Hernquist galactic halo

### v1

@@ I Introduction
Black holes surrounded by galactic dark-matter halos constitute a realistic astrophysical setting where strong gravity and large-scale structure coexist. Observations indicate that both stellar-mass and supermassive black holes reside within extended distributions of dark matter that dominate the galactic mass budget. Although the local dark-matter density near the event horizon is minute compared to the black-hole energy density, even a weakly interacting halo can subtly influence the propagation of perturbative fields and thus modify observable quantities such as quasinormal spectra and Hawking emission rates Liu et al. (2024a); Chen et al. (2024); Konoplya et al. (2025a); Hamil et al. (2025); Pezzella et al. (2025); Bécar et al. (2024); Liu et al. (2024b); Jha (2025); Liu et al. (2025); Mollicone and Destounis (2025); Dubinsky (2025a). Therefore, the study of black-hole perturbations in dark-matter environments has become an important extension of classical perturbation theory, linking gravitational-wave phenomenology with the physics of the surrounding medium.
@@
where $\mu_{\rm eff}$ represents the coupling strength to external matter or fields. Depending on the underlying physics, $\mu_{\rm eff}$ may correspond to the mass of a dark-sector particle, to an effective parameter emerging from self-interactions, or to external influences such as magnetic fields or plasma dispersion Konoplya (2008); Chen et al. (2013); Konoplya and Fontana (2008); Kokkotas et al. (2011); Davlataliev et al. (2025). The introduction of this term alters the asymptotic structure of the effective potential, often producing quasi-bound states or long-lived modes known as quasi-resonances Ohashi and Sakagami (2004); Konoplya and Zhidenko (2018). These long-lived oscillations can substantially affect the late-time tails of perturbations and modify the ringdown signal of gravitational waves Jing (2005); Rogatko and Szyplowska (2007); Koyama and Tomimatsu (2001, 2002); Moderski and Rogatko (2001); Churilova et al. (2020); Dubinsky (2024a). In addition, massive fields may play a role in ultra-low-frequency radiation accessible to Pulsar Timing Arrays Konoplya and Zhidenko (2024a), and appear naturally as effective degrees of freedom in brane-world scenarios, where the bulk-induced tidal term acts as a mass contribution Seahra et al. (2005); Ishihara et al. (2008).
@@
From the theoretical standpoint, the presence of a massive field provides a mechanism for energy exchange between the black hole and its surroundings. If the effective potential allows partial wave trapping, the system may develop superradiant instabilities, whose onset depends on both the mass of the field and the parameters describing the dark-matter halo. Consequently, analyzing the dependence of QNMs on these parameters provides insight into the dynamical stability of the configuration and into possible observational signatures of dark matter or external fields. Detailed studies of perturbations and QNMs of massive fields in various modified or dark-matter–inspired geometries can be found in Lütfüoğlu (2025a); Skvortsova (2025a); Zinhailo (2024a); Konoplya et al. (2025b); Bolokhov (2024a); Lütfüoğlu (2025b); Malik (2025a); Skvortsova (2025b); Bolokhov (2024b); Fernandes et al. (2022); Lütfüoğlu (2025c); Konoplya and Zhidenko (2007); Bolokhov (2024c); Zinhailo (2024b); Dubinsky and Zinhailo (2024); Lütfüoğlu (2025d); Bolokhov (2025); Percival and Dolan (2020).
@@
Beyond the dynamical response, another key characteristic of black-hole perturbations is encoded in the grey-body factors (GBFs), which describe the probability that Hawking radiation quanta escape to infinity instead of being reflected by the effective potential barrier. GBFs determine the energy and particle emission rates and thus directly influence the evaporation process Page (1976a, b); Kanti and March-Russell (2002). Within the context of gauge/gravity duality, they also have a holographic interpretation: the poles of the transmission coefficient correspond to those of retarded Green’s functions in the dual field theory, linking black-hole scattering to transport properties such as optical conductivity in strongly coupled plasmas or holographic superconductors Horowitz and Roberts (2008); Herzog (2009); Konoplya and Zhidenko (2010).
@@
In Konoplya (2021); Cardoso et al. (2022) QNMs and GBFs were found for massless fields in the background of the black hole immersed in a dark-matter halo described by the Hernquist halo Hernquist (1990). Here, we extend that study to the analysis of the spectrum and GBFs for a massive field, where the effective mass term could be interpreted as a contribution from the surrounding magnetic field. The analysis of the spectral and scattering properties of massive fields in the black hole against variations of the surrounding dark-matter environment reveals richer physics in both the time and frequency domains.
@@ II.1 Black hole solution
Following the analysis developed in Cardoso et al. (2022), the gravitational background generated by a galactic halo can be effectively described by the Hernquist density profile Hernquist (1990), which has been successfully employed to reproduce the observed Sérsic surface-brightness distributions of elliptical galaxies and galactic bulges. The matter density is modeled as
@@
When this matter distribution is used as the source of the Einstein equations, the temporal metric component $f(r)$ obtained in Cardoso et al. (2022) takes the form
@@
QNMs and GBFs of massless fields for this solution have been analyzed in Konoplya (2021), while various generalizations for the background metric for more general profiles of halo were found in Konoplya and Zhidenko (2022).
@@ II.2 Perturbations of a massive scalar field
When the mass of the field becomes sufficiently large, the potential peak gradually becomes smoother and may eventually disappear (see Figs. 1 and 2). Notice that for unrealistically dense halos with $M\sim a_{0}$, the effective potential of even massless perturbations develops a double–peak structure (see, for instance, Fig. 3). As a result, strong echoes are expected to appear. Since such echoes are clearly not observed, this provides an additional reason for excluding the regime of exceptionally dense halos from our analysis.
@@ III.1 Time-domain evolution and frequency extraction
The signal $\Psi(t,r)$ recorded at a fixed radius exhibits three distinct regimes: an initial transient depending on the chosen initial data, an exponentially damped ringing governed by the QNMs, and a late-time power-law or exponential tail. To extract the complex frequencies from the intermediate regime, we apply the Prony method Konoplya and Zhidenko (2011), which assumes that the waveform can be approximated as a finite sum of exponentially damped oscillations:
@@
The time-domain evolution is especially valuable when the effective potential develops multiple extrema or when quasi-bound states appear for massive fields, since, in such cases, frequency-domain approaches may converge poorly. In addition, the numerical signal allows one to visualize the entire relaxation process and to assess the stability of the background spacetime against small perturbations. Extensive applications of this approach to a variety of black-hole configurations can be found in Malik (2025b); Konoplya and Zhidenko (2014); Dubinsky (2024b); Skvortsova (2024a); Konoplya et al. (2020); Malik (2025c); Momennia (2022); Aneesh et al. (2018); Malik (2024); Qian et al. (2022); Konoplya and Stashko (2025); Churilova et al. (2021); Dubinsky (2024c, 2025b); Malik (2025d); Cuyubamba et al. (2016); Stuchlík and Zhidenko (2025).
@@ III.2 WKB approach and Padé resummation
To complement the time-domain analysis and obtain semi-analytic estimates of the spectrum, we also employ the higher-order WKB approximation combined with Padé rational expansions. The WKB technique provides an efficient way to determine both quasinormal frequencies and transmission coefficients for smooth, single-barrier potentials of the type encountered in spherically symmetric black-hole perturbations Lütfüoğlu (2025e); Skvortsova (2024b); Bolokhov and Ivashchuk (2022); del Corral and Olmedo (2022); Konoplya (2002); Lütfüoğlu et al. (2025a); Skvortsova (2024c); Momennia (2022); Zhao et al. (2022); Skvortsova (2024d); Kodama et al. (2010); Lütfüoğlu et al. (2025b); Bolokhov and Skvortsova (2025b).
@@ IV Quasinormal modes
The overall behavior of the QNM spectra implies that, in the Hernquist halo, the presence of a massive scalar field can generate long-lived oscillations. The gradual reduction of the damping rate with increasing field mass may enhance the contribution of such modes to the late-time tails of perturbations, while the dependence on the halo scale parameter ensures that these effects remain local to the black-hole vicinity. Therefore, the obtained results support the robustness of black-hole quasinormal ringing against realistic dark-matter environments and clarify how the mass of the perturbed field, rather than the large-scale halo properties, governs the lifetime of the modes. These features are in agreement with observations for the other models of the galactic halo and for massless fields  Konoplya (2021); Konoplya and Zhidenko (2022). Altogether, this brings us to the conclusion that observations of gravitational waves are safe against possible contamination by the galactic halo environment, unless the halo is exceptionally dense in the central region.
@@ V Grey-body factors and absorption cross-sections
with $V_{0}$ and $V_{0}^{\prime\prime}$ denoting the value and the second derivative of the effective potential at its maximum. This expression provides a reliable approximation of the true GBF $\Gamma(\omega)$ along the real frequency axis and has been widely employed in the literature for analyzing black-hole radiation and transmission properties Konoplya et al. (2023); Dubinsky (2025c, d, e); Lütfüoğlu (2025b); Lütfüoğlu et al. (2025b); Lütfüoğlu (2025f); Bolokhov and Skvortsova (2025c).
The relation between the transmission coefficient and the QNM spectrum was first identified in Konoplya and Zhidenko (2024b) and later examined and refined in a number of works Bolokhov and Skvortsova (2025d); Skvortsova (2025c); Malik (2025e, f, g, h). This so-called GBF/QNM correspondence reveals that the poles of the transmission coefficient in the complex-frequency plane coincide with the quasinormal frequencies, at least in the eikonal regime where the WKB approximation is most accurate. At finite multipole numbers $\ell$, deviations appear, and the correspondence becomes approximate—as illustrated by the data obtained in the present study.
The correspondence fails in the same cases where the traditional relation between QNMs and null geodesics breaks down. This occurs, for instance, in higher-curvature theories of gravity where the effective potential at large $\ell$ no longer assumes the universal centrifugal form $V(r)\sim g_{tt}\,\ell(\ell+1)/r^{2}$ Konoplya and Stuchlík (2017); Bolokhov (2024d); Konoplya and Zhidenko (2017). A partial loss of correspondence is also known for asymptotically de Sitter geometries Konoplya (2023), where the presence of a cosmological horizon alters the boundary conditions of the scattering problem.
A more refined relation valid for black holes and wormholes—including the leading and next-to-leading corrections beyond the eikonal regime—was derived in Konoplya and Zhidenko (2024b); Bolokhov and Skvortsova (2025d). In this formulation, the imaginary part of the function $\mathcal{K}$, which governs the transmission probability, can be expressed as Konoplya and Zhidenko (2024b); Bolokhov and Skvortsova (2025d)
@@
where $\Gamma_{\ell}(\omega)$ is the grey-body (transmission) coefficient for the corresponding multipole. In the low-frequency limit, $\sigma_{\mathrm{abs}}$ approaches the area of the event horizon, $\sigma_{\mathrm{abs}}\to A_{\mathrm{H}}$, while at high frequencies it oscillates around the geometric-optics capture cross-section associated with the unstable photon orbit. Thus, $\sigma_{\mathrm{abs}}(\omega)$ provides a bridge between the wave and particle descriptions of radiation, encapsulating both quantum and classical aspects of black-hole interaction with external fields.

### v2

@@ I Introduction
Black holes surrounded by galactic dark-matter halos constitute a realistic astrophysical setting where strong gravity and large-scale structure coexist. Observations indicate that both stellar-mass and supermassive black holes reside within extended distributions of dark matter that dominate the galactic mass budget. Although the local dark-matter density near the event horizon is minute compared to the black-hole energy density, even a weakly interacting halo can subtly influence the propagation of perturbative fields and thus modify observable quantities such as quasinormal spectra and Hawking emission rates. In addition, the spectral and other types of instabilities have been extensively discussed in this context. The recent literature on the above topics is extensive and we mention here only some of these works  Liu et al. (2024a); Chen et al. (2024); Konoplya et al. (2025a); Hamil et al. (2025); Pezzella et al. (2025); Konoplya and Zhidenko (2024a); Bécar et al. (2024); Liu et al. (2024b); Konoplya (2023a); Jha (2025); Liu et al. (2025); Mollicone and Destounis (2025); Dubinsky (2025a); Cheung et al. (2022); Destounis et al. (2025); Rosato et al. (2024); Courty et al. (2023); Mascher et al. (2022). Therefore, the study of black-hole perturbations in dark-matter environments has become an important extension of classical perturbation theory, linking gravitational-wave phenomenology with the physics of the surrounding medium Berti et al. (2025); Destounis and Duque (2023).
@@
where $\mu_{\rm eff}$ represents the mass of the field. Depending on the underlying physics, $\mu_{\rm eff}$ may correspond to either the mass of a field propagating in the black hole’s background or an effective mass corresponding, for example, to the background magnetic field Konoplya (2008); Chen et al. (2013); Konoplya and Fontana (2008); Kokkotas et al. (2011); Davlataliev et al. (2025). The introduction of this term alters the asymptotic structure of the effective potential, often producing quasi-bound states or long-lived modes known as quasi-resonances Ohashi and Sakagami (2004); Konoplya and Zhidenko (2018). These long-lived oscillations can substantially affect not only the ringdown frequencies but also the time-domain evolution of perturbations and modify the ringdown signal of gravitational waves. In addition, the quasi-bound states take place when the mass of the field is turned on. Extensive literature on the above questions can be found in Jing (2005); Rogatko and Szyplowska (2007); Konoplya et al. (2007, 2018); Koyama and Tomimatsu (2001, 2002); Konoplya (2002a); Moderski and Rogatko (2001); Konoplya (2006); Churilova et al. (2020); Dubinsky (2024a); Destounis (2019); Vlachos et al. (2021); Chatzifotis et al. (2022); Vieira et al. (2022); Steinhauer et al. (2025); Vieira and Destounis (2025); Vieira et al. (2025, 2023); Dubinsky (2025b); Malik (2025a); Konoplya and Zhidenko (2006). In addition, massive fields may play a role in ultra-low-frequency radiation accessible to Pulsar Timing Arrays Konoplya and Zhidenko (2024b), and appear naturally as effective degrees of freedom in brane-world scenarios, where the bulk-induced tidal term acts as a mass contribution Seahra et al. (2005); Ishihara et al. (2008).
@@
From the theoretical standpoint, the presence of a massive field provides a mechanism for energy exchange between the black hole and its surroundings. If the effective potential allows partial wave trapping, the system may develop superradiance and superradiant instabilities Starobinskil and Churilov (1974); Starobinskii (1973); Brito et al. (2015), whose onset depends on both the mass of the field and the parameters describing the dark-matter halo. Consequently, analyzing the dependence of QNMs on these parameters provides insight into the dynamical stability of the configuration and into possible observational signatures of dark matter or external fields. Detailed studies of perturbations and QNMs of massive fields in various modified or dark-matter–inspired geometries can be found in Lütfüoğlu (2025a); Skvortsova (2025a); Zinhailo (2024a); Konoplya et al. (2025b); Bolokhov (2024a); Lütfüoğlu (2025b); Malik (2025b); Skvortsova (2025b); Bolokhov (2024b); Fernandes et al. (2022); Lütfüoğlu (2025c); Konoplya and Zhidenko (2007); Bolokhov (2024c); Zinhailo (2024b); Dubinsky and Zinhailo (2024); Konoplya (2024); Lütfüoğlu (2025d); Bolokhov (2025); Percival and Dolan (2020).
@@
Beyond the dynamical response, another key characteristic of black-hole perturbations is encoded in the grey-body factors (GBFs), which describe the probability that Hawking radiation quanta escape to infinity instead of being reflected by the effective potential barrier. GBFs determine the energy and particle emission rates and thus directly influence the evaporation process Page (1976a, b); Kanti and March-Russell (2002). Interestingly, GBFs are much more stable to deformations of the background, they can be used in order to describe the ringdown signal, and even construct the signal through destabilized QNM spectra that appear from modified black holes by astrophysical environments Oshita (2024); Okabayashi and Oshita (2024); Oshita et al. (2024, 2025); Oshita and Cardoso (2025); Rosato et al. (2024); Konoplya and Pappas (2025). Within the context of gauge/gravity duality, they also have a holographic interpretation: the poles of the transmission coefficient correspond to those of retarded Green’s functions in the dual field theory, linking black-hole scattering to transport properties such as optical conductivity in strongly coupled plasmas or holographic superconductors Horowitz and Roberts (2008); Herzog (2009); Konoplya and Zhidenko (2010).
@@
In Konoplya (2021); Cardoso et al. (2022a) QNMs and GBFs were found for massless fields in the background of the black hole immersed in a dark-matter halo described by the Hernquist halo Hernquist (1990). Here, we extend that study to the analysis of the spectrum and GBFs for a massive field, where the effective mass term could be interpreted as a contribution from the surrounding magnetic field. The analysis of the spectral and scattering properties of massive fields in the black hole against variations of the surrounding dark-matter environment reveals richer physics in both the time and frequency domains.
@@ II.1 Black hole solution
Following the analysis developed in Cardoso et al. (2022a), the gravitational background generated by a galactic halo can be effectively described by the Hernquist density profile Hernquist (1990), which has been successfully employed to reproduce the observed Sérsic surface-brightness distributions of elliptical galaxies and galactic bulges. The matter density is modeled as
@@
When this matter distribution is used as the source of the Einstein equations, the temporal metric component $f(r)$ obtained in Cardoso et al. (2022a) takes the form
@@
QNMs and GBFs of massless fields for this solution have been analyzed in Konoplya (2021), while various generalizations for the background metric for more general profiles of halo were found and studied in Konoplya and Zhidenko (2022); Figueiredo et al. (2023); Speeney et al. (2024); Pezzella et al. (2025); Fernandes and Cardoso (2025); Destounis and Fernandes (2025); Jusufi (2023); Konoplya and Zhidenko (2025).
@@ II.2 Perturbations of a massive scalar field
When the mass of the field becomes sufficiently large, the potential peak gradually becomes smoother and may eventually disappear (see Figs. 1 and 2). Notice that for unrealistically dense halos with $M\sim a_{0}$, but still corresponding to a black hole solution, the effective potential of even massless perturbations develops a double–peak structure (see, for instance, Fig. 3). As a result, strong echoes are expected to appear. Since such echoes are clearly not observed, this provides an additional reason for excluding the regime of exceptionally dense halos from our analysis.
@@ III.1 Time-domain evolution and frequency extraction
The signal $\Psi(t,r)$ recorded at a fixed radius exhibits three distinct regimes: an initial transient depending on the chosen initial data, an exponentially damped ringing governed by the QNMs, and a late-time power-law or exponential tail. The quasinormal ringing stage is the superposition of the dominant fundamental mode and an infinite tower of overtones. To extract the complex frequencies from the intermediate regime, we apply the Prony method Berti et al. (2006, 2007); Konoplya and Zhidenko (2011), which assumes that the waveform can be approximated as a finite sum of exponentially damped oscillations:
@@
The time-domain evolution is especially valuable when the effective potential develops multiple extrema or when quasi-bound states appear for massive fields, since, in such cases, frequency-domain approaches may converge poorly. In addition, the numerical signal allows one to visualize the entire relaxation process and to assess the stability of the background spacetime against small perturbations. Extensive applications of this approach to a variety of black-hole configurations can be found in Malik (2025c); Konoplya and Zhidenko (2014); Dubinsky (2024b); Skvortsova (2024a); Konoplya et al. (2020); Malik (2025d); Momennia (2022); Aneesh et al. (2018); Malik (2024); Qian et al. (2022); Konoplya and Stashko (2025); Churilova et al. (2021); Dubinsky (2024c, 2025c); Malik (2025e); Cuyubamba et al. (2016); Stuchlík and Zhidenko (2025).
@@ III.2 WKB approach and Padé resummation
To complement the time-domain analysis and obtain semi-analytic estimates of the spectrum, we also employ the higher-order WKB approximation combined with Padé rational expansions. The WKB technique provides an efficient way to determine both quasinormal frequencies and transmission coefficients for smooth, single-barrier potentials of the type encountered in spherically symmetric black-hole perturbations Lütfüoğlu (2025e); Skvortsova (2024b); Bolokhov and Ivashchuk (2022); del Corral and Olmedo (2022); Konoplya (2002b); Lütfüoğlu et al. (2025a); Skvortsova (2024c); Momennia (2022); Zhao et al. (2022); Skvortsova (2024d); Kodama et al. (2010); Lütfüoğlu et al. (2025b); Bolokhov and Skvortsova (2025b).
@@ IV Quasinormal modes
The overall behavior of the QNM spectra implies that, in the Hernquist halo, the presence of a massive scalar field can generate long-lived oscillations. The gradual reduction of the damping rate with increasing field mass may enhance the contribution of such modes to the late-time tails of perturbations, while the dependence on the halo scale parameter ensures that these effects remain local to the black-hole vicinity. Therefore, the obtained results support the robustness of black-hole quasinormal ringing against realistic dark-matter environments and clarify how the mass of the perturbed field, rather than the large-scale halo properties, governs the lifetime of the modes. These features are in agreement with observations for the other models of the galactic halo and for massless fields  Konoplya (2021); Konoplya and Zhidenko (2022). Altogether, this brings us to the conclusion that possible observations of minimally coupled waves must be safe against possible contamination by the galactic halo environment, unless the halo is exceptionally dense in the central region. However, the case of coupled fields deserves separate consideration Cardoso et al. (2022b).
@@ V Grey-body factors and absorption cross-sections
with $V_{0}$ and $V_{0}^{\prime\prime}$ denoting the value and the second derivative of the effective potential at its maximum. This expression provides a reliable approximation of the true GBF $\Gamma(\omega)$ along the real frequency axis and has been widely employed in the literature for analyzing black-hole radiation and transmission properties Konoplya et al. (2023); Dubinsky (2025d, 2026, e); Lütfüoğlu (2025b); Lütfüoğlu et al. (2025b); Lütfüoğlu (2025f); Bolokhov and Skvortsova (2025c).
The relation between the transmission coefficient and the QNM spectrum was first identified in Konoplya and Zhidenko (2024c) and later examined and refined in a number of works Bolokhov and Skvortsova (2025d); Skvortsova (2025c); Malik (2025f, g, h, i). This so-called GBF/QNM correspondence reveals that the poles of the transmission coefficient in the complex-frequency plane coincide with the quasinormal frequencies, at least in the eikonal regime where the WKB approximation is most accurate. At finite multipole numbers $\ell$, deviations appear, and the correspondence becomes approximate—as illustrated by the data obtained in the present study.
The correspondence fails in the same cases where the traditional relation between QNMs and null geodesics breaks down. This occurs, for instance, in higher-curvature theories of gravity where the effective potential at large $\ell$ no longer assumes the universal centrifugal form $V(r)\sim g_{tt}\,\ell(\ell+1)/r^{2}$ Konoplya and Stuchlík (2017); Bolokhov (2024d); Konoplya and Zhidenko (2017). A partial loss of correspondence is also known for asymptotically de Sitter geometries Konoplya (2023b), where the presence of a cosmological horizon alters the boundary conditions of the scattering problem.
A more refined relation valid for black holes and wormholes—including the leading and next-to-leading corrections beyond the eikonal regime—was derived in Konoplya and Zhidenko (2024c); Bolokhov and Skvortsova (2025d). In this formulation, the imaginary part of the function $\mathcal{K}$, which governs the transmission probability, can be expressed as Konoplya and Zhidenko (2024c); Bolokhov and Skvortsova (2025d)
@@
where $\Gamma_{\ell}(\omega)$ is the grey-body (transmission) coefficient for the corresponding multipole. In the low-frequency limit, $\sigma_{\mathrm{abs}}$ approaches the area of the event horizon, $\sigma_{\mathrm{abs}}\to A_{\mathrm{H}}$, while at high frequencies it oscillates around the geometric-optics capture cross-section associated with the unstable photon orbit. Thus, $\sigma_{\mathrm{abs}}(\omega)$ provides a bridge between the wave and particle descriptions of radiation, encapsulating both quantum and classical aspects of black-hole interaction with external fields Futterman et al. (2012).